OpenStudy (anonymous):

I NEED SOME HELP PLEASE A group of college students built a self-guided rover and tested it on a plane surface. They programmed the vehicle to move along the path A–B–C–D–A represented on the coordinate plane. What distance will the rover cover if it completes one circuit? 36 meters 40 meters 54 meters 60 meters 68 meters

OpenStudy (anonymous):

Where is the graph image? @asela121

OpenStudy (anonymous):

thank you but i answered the question already my teacher helped me... do you think you can help me with another problem?

OpenStudy (anonymous):

@bajji

OpenStudy (anonymous):

ya... if i can... :)

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

If A(x1, y1), B(x2, y2), C(x3, y3), and D(x4, y4) form two line segments, AB and CD , which condition needs to be met to prove AB and CD? the image can help

OpenStudy (anonymous):

It says the condition is AB perpendicular CD so, the (Slope of AB)*(Slope of CD)=-1 If the slope AB is m, then the slope of a CD(say, m2) would be -1/m this means that m2 = -1/m and so m2 x m = -1 Now can you find which is the right option? :) @asela121

OpenStudy (anonymous):

yes :) so it would be the third one right ? @bajji

OpenStudy (anonymous):

Exactly! :)

OpenStudy (anonymous):

thank you :D